# thanks: The National Key Research and Development Program of China (Project No. 2022YFA1602001)

### v1

@@ Abstract
For tracking and tagging the cosmic-ray muon (CR-muon), the Taishan Antineutrino Observatory (TAO) experiment is equipped with a top veto tracker (TVT) system composed of 160 modules, each consisting of plastic scintillator (PS) strip as target material, embedded wavelength shifting fiber (WLS-fiber) as photon collection and transmission medium, and silicon photomultipliers (SiPMs) at both ends as read-out. This article introduces the unique design of the module and reports the excellent performance of all modules, providing guidance and important reference for the process design of scintillation detectors with WLS-fibers. In general, when the CR-muon hits the center of plastic scintillator and without optical grease, the most probable value of the signal amplitude at one end of the PS strip is greater than 40.8 p.e. and 51.5 p.e. for all the 2 m-length modules and 1.5 m-length modules respectively. The CR-muon tagging efficiency of PS module is measured to be more than 99.3%, which meets the requirement of TAO.
Plastic scintillator, WLS-fiber, Muon tagging efficiency, Light yield, JUNO-TAO
@@ I Introduction
For most neutrino or low background detectors, especially those near the ground Hakenmuller:2019ecb ; TAO-CDR ; NEON:2022hbk ; Singh:2017jow ; Ang:2019ard , a muon veto system is necessary to tag muon-induced particles going to the main detector NUCLEUS:2022vyj ; JUNO:2023cbw ; Li:2022wqc ; Seo:2022vzr . Cosmic ray (CR) muons can induce neutrons, producing gammas that mimic the coincidence signals. Therefore, efficient identification of muons to remove associated events is essential in the such experiments, which is the primary goal of the muon veto detector JUNO:2023cbw ; Rico:2022aoc ; Coveyou:2023fja ; Adam:2007ex . Plastic scintillation has been adopted as the basic unit of anti-coincidence detectors in many experiments, due to its advantages of easy machining, flexible structure design, stable performance, and good adaptability Moiseev:2007zz ; Perotti:2006yc ; Hao-Ran2021Discrimination ; Zhuo2021Adaptability .
Taishan Antineutrino Observatory (TAO) TAO-CDR , scheduled for running in end of 2024, independently measures the antineutrino energy spectrum of the reactor with unprecedented energy resolution for the Jiangmen Underground Neutrino Observatory (JUNO) TAO-CDR ; JUNO-CDR ; JUNO:2015zny ; JUNO:2021vlw . TAO will provide a unique reference for other experiments and nuclear databases Capozzi:2020cxm . A Top Veto Tracker (TVT) is designed serving as a part of the muon veto system, where the Plastic Scintillator (PS) module is the key element of the TVT. A detailed description of the TVT is given in Ref Luo:2023inu , where most aspects of the system design and optimizing process are covered.
This paper aims to describe the unique designs and superior performances. We will introduce the PS module assembly and the production in sec. II. In sec. III, the testing system and the measurement results of one module will be presented from three main aspects: the light yield along its length of the PS module by the measurement of the CR muon, effective attenuation length, and coupling effect between wavelength shifting fiber (WLS-fiber) and silicon photomultipliers (SiPMs). In sec. IV, details of the module performance will be discussed. In sec. V, the detection efficiency of the module is demonstrated under different threshold conditions. Finally, a summary is given in sec. VI.
@@ II Module production
Both 2000mm-PS and 1500mm-PS are fabricated by Beijing Hoton Nuclear Technology Co., Ltd.Haotang:2014 . The PS modules are made with extruded plastic scintillator type of SP101 polymerized with liquid polystyrene added with P-triphenyl and POPOP. The WLS-fiber BCF92 Tur:2009en ; DietzLaursonn:2016tpy with a diameter of 1.5 mm is used and its end surface is flat. Two fibers are focused into a single group that can be coupled with optical sensors.
@@
As shown in Fig. 3, the brightness of optical fiber and the PS can still be distinguished even without shading treatment. SiPM windows of dimension $4\times 4mm^{2}$ have been cut on the backend of the covered PS. Fig. 3 shows the image of one end of the PS module, eight bright spots in four groups can be clearly seen from the $4\times 4mm^{2}$ optical windows, which are the positions of the optical fibers.
@@
Four SiPMs of J-series MicroJ-40035-TSV citeonsemi with integrated readout are used as the photon sensor at each end of the PS module. Fig. 4 shows the layout of SiPMs on a PCB and the circuit of the output signals channel. Four SiPMs (SiPM1, SiPM2, SiPM3, and SiPM4) are paired the optical windows, the designed PCB can read-out each SiPM signal and also summed the four SiPM signals. According to previous simulation results from Ref Luo:2023inu , the difference among the four SiPMs on one end are very small. The design of reading out each SiPM signal is precisely to measure and verify the difference among the four SiPM signals. The design of four SiPM summing channels is to reduce the number of readout channels, and the final design of the PCB will only have one summing channel output. At present, the design of this type of PCB is mainly for module quality inspection and performance testing. The current version of PCB are designed and fabricated by Dual Rainbow Technology Co., Ltd.dualrainbow . Fig. 4 is the photo of the backend of the PS module during mass testing after the PCB mounted. The aluminum fixtures are used to mount the PCB onto the plastic scintillator. After the installation of the fixture, they are packaged as a whole and finally coated with black epoxy for light-tight treatment. The output signal line uses coaxial cables with SMA connectors.
@@ III Testing system and measurements
Figure 5: A side view of the testing system and a flowchart of the data flow are used to measure the performance of each PS strip by muon at 9 locations.
@@
Fig. 5 demonstrates the components of the testing system. The long strip placed horizontally on the platform is the 2000mm-PS. A CR muon monitor system consists of two plastic scintillators with the type of EJ200 cite:Ej200 with dimensions of $6\times 6\times 18\,cm^{3}$ and a 2-inch Photo Multiplier Tube (PMTs) of type XP3232 cite:Hpmt . The coincidence of the two monitors is used as a trigger to identify a muon passing through the 2000mm-PS, and to characterise the performance of the strip. The vertical distance between the two monitors is 24 cm to tag muons going through the 2000mm-PS strip. In Fig. 5, the two CR monitors are placed in the center of the PS module. A high-resolution oscilloscope with type lecroy-HDO4104A cite:LecOsc is used to obtain a total of four channels of signals, including the two of the CR monitors, another two of the summation or single channels of the SiPMs attached to the two ends of the PS strip. In addition, another fast high-resolution time an amplitude digitizer with type MDPP-32-QDC cite:Mesytec with less dead time is also used in parallel for comparison, which is internally realized as a 32-channel adjustable low noise amplifier and a variable differentiation stage, followed by band pass filters and 80 MHz sampling ADCs. MDPP-32-QDC can realize about eight times more channels than the used oscilloscope. For subsequent measurements, we can perform 8 PS strips measurements in parallel, and simultaneously measure the single and summed channels at the same time.
@@
Figure 6: (a) the amplitude spectrum of SiPM1. (b) the amplitude spectrum of SiPM2. (c) the amplitude spectrum of the summation channels of one end. (d) the amplitude spectrum of the summation channels of the other end. (e) The distribution map of the light yield of a 2000mm-PS strip with the CR monitors at different locations. (f) MDPP-32-QDC data vs. Oscilloscope data.
@@
In the test, we warmed up the PMTs of the CR monitors and set their operating voltages both to -1020 V. With a pre-test and calibration, the threshold for both the PMTs of the CR monitors is set to -14 mV for muon identification.
@@
Fig. 6 shows the distribution of the amplitude of the signal from SiPM1. Fig. 6 is the distribution of the amplitude of the signal from SiPM2 from the same PCB. The black dots with error bars are the experimental data. The spectrum is fitted by a joint function of exponential and Landau distribution for gamma and muons. For SiPM1, the most probable value (MPV) of signal amplitude for muon is 119.5 mV. For SiPM2, the MPV of signal amplitude for muon is 122.1 mV. The difference between the SiPM1 and SiPM2 is < 3 mV, which is consistent with the previous simulation Luo:2023inu , where the signals from different optical fibers are relatively equal in strength.
Fig. 6 shows the distribution of the signal amplitude from the summation channel of one end of the PS when the CR monitors located at its center, where the black dots with error bars are experimental data, and he red line is the image of the fitting function. The MPV of the fitted Landau distribution is 465.6 mV. Fig. 6 is the distribution of another end of the same PS. The MPV of the fitted Landau distribution is 467.8 mV, which is almost four times of an individual SiPM channel as in Fig. 6.
@@
To characterize the signal strength when the muon going through the plastic scintillator strips, the MPV fitted by Landau distribution can be used as an estimation of the light yield of the strip after converted into p.e., which will be measured and calculated at nine different locations along thelength direction of the PS module. Fig. 6 is measured results, where the x-axis is the position of the CR monitor in the longitudinal direction of the PS strip, and the error bar in x-axis represents the 6 cm size width of the CR monitor. The y-axis is the measured light yield in p.e. units, which is calculated by converting the relationship between MPV and the signal amplitude of a single photoelectron with a pre-calibrated factor (10 mV per p.e., claimed efficiency 38% under the bias voltage 3.5 V, cross talk ration 11.6%). The black, red, and green lines represent the light yield of the 4 SiPM summation channel of one end, the other end and the sum of both ends, respectively. When the CR monitor is located in the range of -80 or 80 cm of the PS strip, the minimum light yield can still be higher than 30 p.e. for single end, and the maximum light yield can almost reach 80 p.e. at single end.
@@
In Fig. 7, by fitting the logarithm of light yield with distance from SiPMYang:2022awv . we obtained an effective attenuation length of 241 ± 6.97 cm and it meet our design requirement of > 200 cm.
For a cross-check, we also compared the results from the two sets of data acquisition systems (oscilloscope or MDPP-32-QDC). Placing the CR monitors at the center of the PS strip, we measured seven PS strips using the two systems, and the MPV of each PS strip was fitted and obtained. In Fig. 6, the x-axis is the MPV from oscilloscope (in p.e.), and the y-axis is the MPV from MDPP-32-QDC (in ADC). The seven blue inverted triangle points are the MPVs of the seven PS strips, and fitted by a linear function. The red line is the fitted results. It is around 11.3 ADC/p.e.
@@ IV Module performance
We conducted four batches of acceptance and performance tests following the production in total. Fig. 8 shows the MPV of single end of all the plastic scintillator strips when the CR monitors are located at the centre of the PS module. It can be seen in Fig. 8, the typical light yield of single end of the 108 strips of 2000mm-PS modules (blue line) is about 40.8 p.e., and the majority measured to be 46 p.e. It can be obtained from the red line for the 52 strips of 1500mm-PS strips, the typical light yield of single end is about 51.5 p.e., and the majority concentrated at 55 p.e.
In addition, we tried EJ-550 cite:Ej550 silicone optical grease for the coupling between SiPM and WLS-fiber, rather than the coupling by only air during the test. Subsequent measurements were conducted to demonstrate that the coupling between SiPM and WLS-fiber by optical grease can increase the measured light yield by at least 5 p.e. at one end (around 12.5% in light yield for 2000mm-PS modules).
@@ V Detection efficiency
$Eff_{one}=\frac{N[one]_{det}}{N_{reg}};$ (2)
$Eff_{other}=\frac{N[other]_{det}}{N_{reg}}$ (3)
$Eff_{module}=\frac{N[module]_{det}}{N_{reg}}$ (4)
$Eff_{oneANDother}=\frac{N[oneANDother]_{det}}{N_{reg}}$ (5)
@@
$Eff_{one}$(%) $99.72\pm 0.03$ $99.69\pm 0.03$ $99.53\pm 0.03$
$Eff_{other}$(%) $99.70\pm 0.03$ $99.62\pm 0.03$ $99.46\pm 0.03$
$Eff_{oneANDother}$(%) $99.63\pm 0.03$ $99.48\pm 0.03$ $99.01\pm 0.04$
@@
$Eff_{module}$(%) $99.70\pm 0.03$ $99.69\pm 0.03$ $99.62\pm 0.03$
Tab. 1 shows the results of calculated detection efficiency with different thresholds on SiPM signal and configurations. From the overall perspective of the Tab. 1, it can be observed that as the threshold of the SiPM signal increases, the detection efficiency gradually decreases as expected. From Tab. 1, it can also be observed that $Eff_{oneANDother}$ decreases the fastest as the threshold increases. Fig. 11 displayed the energy spectrum under "AND" logic model. It can be observed that in this mode, there is a significant decrease in the coincidence background.
@@
Tab. 2 shows the rate vs. threshold of the sum of both ends of the PS module, and the coincidence of the two ends when the PS module located at the laboratory on ground. From Tab. 2, it can also be observed that under this "AND" logic mode, the events of backgrounds that coincidentally match sharply decreases.
@@
The CR monitors were also placed at other positions of the PS modules to check the efficiency. When the CR monitors are off to the center of the PS module, the signal strength obtained at the two ends of the PS module will be different. Fig. 12 shows the 2-D plot of the two ends of the PS module when the muon hits around 60 cm away from the center of the PS module. It can be clearly seen that the noise or environmental background can be identified from the signal of the muon. But the signal at one end is significantly larger than the other end, which could generate a difference in the efficiency of the two ends with the same threshold on SiPM signals. Considering the possible efficiency and background under different thresholds, we ultimately decided to use $Eff_{module}$ and $Eff_{oneANDother}$ as the characterization of overall detection efficiency of PS module.
@@
Fig. 13 shows the relationship between detection efficiency of the sum of the two ends and muon position (CR monitors’ position). The position measurements were made by moving the monitors along the length by a 20 cm step for the 2000mm-PS module (indicated in the figure by dotted lines), or by a 15 cm step for the 1500mm-PS module (indicated in the figure by solid lines). The error bar of the X represents the 6 cm size width of the CR monitor, the Y-coordinate is the detection efficiency of the PS module when the threshold of the sum of the two ends is set to 30 p.e. or 50 p.e., the error bar of the Y-axis is the error in efficiency. As the threshold increases, the detection efficiency corresponding to all positions of the PS module decreases. For a 1500mm-PS module, when the threshold is set to 30 p.e., the detection efficiency of each point is higher than 99.7%. For a 2000mm-PS module, when the threshold is set to 50 p.e., the detection efficiency of each point is higher than 99.6%. Overall, the 1500-mm PS detection efficiency is higher than that of 2000-mm PS. Fig. 13 shows the relationship between detection efficiency of "AND" logic of the two ends and muon position. In this mode, the efficiency of 1500mm-PS is slightly higher than that of 2000mm-PS. For all PS module, when the threshold is set to 20 p.e., the detection efficiency of each point is higher than 99.3%.

### v2

# Performance of plastic scintillator modules for top veto tracker at Taishan Antineutrino Observatory

@@ Abstract
The Taishan Antineutrino Observatory (TAO) experiment incorporates a top veto tracker (TVT) system comprising 160 modules, each composed of plastic scintillator (PS) strips, embedded wavelength shifting fibers (WLS-fibers), and silicon photomultipliers (SiPMs). This article highlights the performance of all produced modules following the production and readout/trigger design, providing insights for scintillation detectors with WLS-fibers. Three kinds of trigger modes and its efficiency have been defined to comprehensively evaluate the performance of this unique design, which has been verified for the batch production, along with comprehensive measurement strategies and quality inspection methods. In "module" mode, the detection(tagging) efficiency of the PS exceeds 99.67% at a 30 photoelectron threshold, and even in "AND" mode, it surpasses 99.60% at a 15 photoelectron threshold. The muon tagging efficiency meets TAO’s requirements. The production and performance of the PS module set a benchmark for other experiments, with optimized optical fiber arrangements that enhance light yield and muon detection efficiency.
Plastic scintillator, WLS-fiber, Muon tagging efficiency, Light yield, TAO
@@ I Introduction
For most neutrino or low background detectors, especially those near the ground Hakenmuller:2019ecb ; TAO-CDR ; SSAFSM ; ZXJCYD ; CFY ; NEON:2022hbk ; Singh:2017jow ; Ang:2019ard , a muon veto system is necessary to tag muon-induced particles going to the main detector NUCLEUS:2022vyj ; JUNO:2023cbw ; Li:2022wqc ; Seo:2022vzr ; W:Chen . Cosmic ray (CR) muons can induce neutrons, producing gammas that mimic the coincidence signals. Therefore, efficient identification of muons to remove associated events is essential in the such experiments, which is the primary goal of the muon veto detector JUNO:2023cbw ; Rico:2022aoc ; Coveyou:2023fja ; Adam:2007ex ; Qain2021 . Plastic scintillation(PS) has been adopted as the basic unit of anti-coincidence detectors in many experimentsWYP2021 , due to its advantages of easy machining, flexible structure design, stable performance, and good adaptability Moiseev:2007zz ; Perotti:2006yc ; Hao-Ran2021Discrimination ; Zhuo2021Adaptability ; ThaK:c ; JND ; YP:Chen ; YT:Qu ; WL2022 ; WuZ2018 ; HP2021 ; ps_ap .
Taishan Antineutrino Observatory (TAO) TAO-CDR ; T:CDR ; TC:DR ; TCD:R ; Zhang:2024znq scheduled to begin operation in spring 2025, will independently measures the antineutrino energy spectrum of the reactor with unprecedented energy resolution for the Jiangmen Underground Neutrino Observatory (JUNO) TAO-CDR ; JUNO-CDR ; JUNO:2015zny ; JUNO:2021vlw ; JUNOcpc . TAO will provide a unique reference for other experiments and nuclear databases Capozzi:2020cxm . A Top Veto Tracker (TVT) is designed serving as a part of the muon veto system, where the PS module is the key element of the TVT. A detailed description of the TVT is given in Ref Luo:2023inu , where most aspects of the system design and optimizing process are covered. In short, it is a design with optimized WLS-fiber arrangement (uniformly bent in PS) that can improve light yield and muon detection efficiency.
This paper aims to describe the unique designs and performances. We will introduce the PS module assembly and the production in sec. II. In sec. III, the testing system and the measurement results of one module will be presented from three main aspects: the light yield along its length of the PS module by the measurement of the CR muon, effective attenuation length, and coupling effect between wavelength shifting fiber (WLS-fiber) and silicon photomultipliers (SiPMs). In sec. IV, details of the module performance will be discussed. In sec. V, the detection efficiency of the module is demonstrated under different threshold conditions. Finally, a summary is given in sec. VI.
@@ II Module production
Both 2000mm-PS and 1500mm-PS are fabricated by Beijing Hoton Nuclear Technology Co., Ltd.. The PS modules are made with extruded plastic scintillator type of SP101 polymerized with liquid polystyrene added with P-triphenyl and POPOP. The WLS-fiber BCF92 Tur:2009en ; DietzLaursonn:2016tpy with a diameter of 1.5 mm is used and its end surface is flat. Two fibers are focused into a single group that can be coupled with optical sensors. To ensure the precise arrangement of WLS-fibers in the design, two methods were utilized: slotting and drilling. Firstly, a slot was made in the middle of the plastic scintillator (4 mm deep), and holes were accurately drilled at both ends of the plastic scintillator. Subsequently, The groove was polished and cleaned. Finally, the WLS-fibers were placed at the bottom of the slot and secured, followed by filling with optical silicone grease (Type: SL600).
@@
As shown in Fig. 3, the brightness of optical fiber and the PS can still be distinguished even without shading treatment. SiPM windows of dimension $4\times 4mm^{2}$ have been cut on the backend of the covered PS. Fig. 3 shows the image of one end of the PS module, eight bright spots in four groups can be clearly seen from the $4\times 4mm^{2}$ optical windows, which are the positions of the WLS-fibers.
@@
Four SiPMs of J-series MicroJ-40035-TSV with integrated readout are used as the photon sensor at each end of the PS moduleLW2021 ; TA2023 ; Sipm_ps ; Si:SiPM . Fig. 4 shows the layout of SiPMs on a PCB and the circuit of the output signals channel. Four SiPMs (SiPM1, SiPM2, SiPM3, and SiPM4) are coupled to the optical windows, the designed PCB can read-out each SiPM signal and also sum up the four SiPM signals. According to previous simulation results from Ref Luo:2023inu , the difference among the four SiPMs on one end are very small. The design of reading out each SiPM signal is to measure and verify precisely the difference among the four SiPM signals. The design of four SiPM summing channels is to reduce the number of readout channels, and the final design of the PCB will only have one summing channel output. At present, the design of this type of PCB is mainly for module quality inspection and performance testing. The current version of PCB are fabricated by Dual Rainbow Technology Co., Ltd.dualrainbow . Fig. 4 is the photo of the backend of the PS module during mass testing after the PCB is mounted. The aluminum fixtures are used to mount the PCB onto the plastic scintillator. After the installation of the fixture, they are packaged as a whole and finally coated with black epoxy for light-tight treatment. The output signal line uses coaxial cables with SMA connectors.
@@ III Testing system and measurements
Figure 5: A flowchart of the data flow are used to measure the performance of each PS strip by muon at 9 locations.
@@
Fig. 5 is a flowchart of the data flow used to measure the performance of each PS strip by muon at 9 locations. A CR muon monitor system consists of two plastic scintillators with the type of EJ200 cite:Ej200 with dimensions of $6\times 6\times 18\,cm^{3}$ and a 2-inch Photo Multiplier Tube (PMTs) Chen2023 ; YF2011 of type XP3232. The coincidence of the two monitors is used as a trigger to identify a muon passing through the 2000mm-PS, and to characterise the performance of the strip. The vertical distance between the two monitors is 24 cm to tag muons going through the 2000mm-PS strip. In Fig. 5, the two CR monitors are placed in the center of the PS module. A high-resolution oscilloscope with type lecroy-HDO4104A is used to obtain a total of four channels of signals, including the two of the CR monitors, and another two of the summation or single channels of the SiPMs attached to the two ends of the PS strip. In addition, another fast high-resolution time and amplitude digitizer with type MDPP-32-QDC with less dead time is also used in parallel for comparison, which is internally realized as a 32-channel adjustable low noise amplifier and a variable differentiation stage, followed by band pass filters and 80 MHz sampling ADCs. MDPP-32-QDC can realize about eight times more channels than the used oscilloscope. For subsequent measurements, we can perform 8 PS strips measurements in parallel, and simultaneously measure the single and summed channels at the same time.
@@
Figure 6: (a) The amplitude spectrum of two adjacent SiPMs at one end. (b) The amplitude spectrum of the summation channels at both ends. (c) The distribution map of the light yield of a 2000mm-PS strip with the CR monitors at different locations. (d) MDPP-32-QDC data vs. Oscilloscope data. (e) The signal spectrum of the PS on one side when the CR monitor is located at -97  cm(the outermost edge). (f) The signal spectrum of the PS on the other side when the CR monitor is located at -97  cm.
@@
In the test, we warmed up the PMTs of the CR monitors and set their operating voltages both to -1020 V. With a pre-test and calibration, the threshold for both PMTs of the CR monitors was set to -14 mV for muon identification.
@@
Fig. 6 shows the distribution of the amplitude of the signal from two adjacent SiPMs on the same PCB at one end of the PS. The blue and red dots represent experimental data from SiPM1 and SiPM2, respectively. Except for a few datas that do not overlap, the two spectrum almost overlap in other areas, The spectrum is fitted by a joint function of exponential and Landau distribution for gamma and muons. For SiPM1, the most probable value (MPV) of signal amplitude for muon is 119.5 mV. For SiPM2, the MPV of signal amplitude for muon is 122.1 mV. The difference between the SiPM1 and SiPM2 is < 3 mV, which is consistent with the previous simulation Luo:2023inu , where the signals from different WLS-fibers are relatively equal at the 5% level. This also indicates that our WLS-fibers layout is effective and fairly uniform. Fig. 6 shows the distribution of the amplitude of the signal from the summation channel at both ends of the same PS when the CR monitors located at its center, the blue and red dots represent experimental data from one end and another end, respectively. It is obvious that these two spectra almost overlap. At one end, the MPV of the fitted Landau distribution is 465.6 mV, at another end, the MPV of the fitted Landau distribution is 467.8 mV, which are same and is almost four times of an individual SiPM channel as in Fig. 6.
@@
To characterize the signal strength when the muon going through the plastic scintillator strips, the MPV fitted by Landau distribution can be used as an estimation of the light yield of the strip after converted into p.e., which will be measured and calculated at nine different locations along the longitudinal direction of the PS module. Fig. 6 is measured results, where the x-axis is the position of the CR monitor in the longitudinal direction of the PS strip, and the error bar in x-axis represents the 6 cm size width of the CR monitor. The y-axis is the measured light yield in p.e. units, which is calculated by converting the relationship between MPV and the signal amplitude of a single photoelectron with a pre-calibrated factor (10 mV per p.e., claimed efficiency 38% under the bias voltage 28 V(the overvoltage is 3.5 V), cross talk ration 11.6%). The black, red, and green lines represent the light yield of the 4 SiPM summation channel of one end, the other end and the sum of both ends, respectively. When the CR monitor is located in the range of -80 or 80 cm of the PS strip, the minimum light yield can still be higher than 35 p.e. for single end, and the maximum light yield can almost reach 80 p.e. at single end.
For a cross-check, we also compared the results from the two sets of data acquisition systems (oscilloscope or MDPP-32-QDC). Placing the CR monitors at the center of the PS strip, we measured seven PS strips using the two systems, and the MPV of each PS strip was fitted and obtained. In Fig. 6, the x-axis is the MPV from oscilloscope (in p.e.), and the y-axis is the MPV from MDPP-32-QDC (in ADC). The seven blue inverted triangle points are the MPVs of the seven PS strips, and fitted by a linear function. The red line is the fitted results. It is around 11.3 ADC/p.e., represents the average ADC count value corresponding to each photoelectron. In addition, we sampled and measured the outermost position, placing the CR monitor near -97 cm from the center of the PS strip. Fig. 6 shows the signal spectrum of one end of the PS that is far away from the CR monitor, the MPV of the fitted Landau distribution was 34.31  p.e., Fig. 6 shows the signal spectrum of the other end of the PS, which is near the CR monitor, the MPV of the fitted Landau distribution was 80.07  p.e., similarly, from the comparison between Fig. 6 and Fig. 6, it can be observed that the larger the MPV, the wider the Landau width. The width of the far end fitting is reduced by almost three times compared to the near end fitting width.
@@
In Fig. 7, by fitting the logarithm of light yield with distance from SiPMYang:2022awv ; YXW2020 ; GL2013 . we obtained an effective attenuation length of 241 ± 6.97 cm and it met our design requirement of > 200 cm. According to simulations and some experiments, the intrinsic attenuation length of PSs is approximately 2 meters at a photon wavelength of 410 nm, while the intrinsic attenuation length of WLS-fibers is approximately 3.8 meters at a photon wavelength of 410 nm. The effective attenuation length is greater than the intrinsic attenuation length of the PS, indicating that the optimized arrangement of fiber insertion is effective. It is known that the diameter of WLS-fibers is very small (1.5mm) compared to PSs, with a total volume of less than 5%, and most photons are still attenuated in PSs. Therefore, the effective attenuation length is still mainly caused by the intrinsic attenuation length of PS.
@@ IV Module performance
We conducted four batches of acceptance and performance tests following the production in total. Fig. 8 shows the MPV of single end of all the plastic scintillator strips when the CR monitors are located at the centre of the PS module. It can be seen in Fig. 8 that the lowest light yield from single end of the 108 strips of 2000mm-PS modules (blue line) is about 40.8 p.e., and the mode is measured to be 46 p.e. It can be obtained from the red line for the 52 strips of 1500mm-PS strips, the lowest light yield of single end is about 51.5 p.e., and the mode concentrates at 55 p.e..
In addition, we tried EJ-550 YY550 silicone optical grease for the coupling between SiPM and WLS-fiber, rather than the coupling by only air during the test. Subsequent measurements were conducted to demonstrate that the coupling between SiPM and WLS-fiber by optical grease can increase the measured light yield by at least 5 p.e. at one end (around 12.5% in light yield for 2000mm-PS modules). This indicates that the light yield of our officially operated PS module will be even higher.
@@ V Detection efficiency
$\epsilon_{one}=\frac{N[one]_{det}}{N_{reg}};$ (2)
$\epsilon_{other}=\frac{N[other]_{det}}{N_{reg}}$ (3)
$\epsilon_{module}=\frac{N[module]_{det}}{N_{reg}}$ (4)
$\epsilon_{oneANDother}=\frac{N[oneANDother]_{det}}{N_{reg}}$ (5)
@@
$\epsilon_{one}$(%) $99.72\pm 0.03$ $99.69\pm 0.03$ $99.53\pm 0.03$
$\epsilon_{other}$(%) $99.70\pm 0.03$ $99.62\pm 0.03$ $99.46\pm 0.03$
$\epsilon_{oneANDother}$(%) $99.63\pm 0.03$ $99.48\pm 0.03$ $99.01\pm 0.04$
@@
$\epsilon_{module}$(%) $99.70\pm 0.03$ $99.69\pm 0.03$ $99.62\pm 0.03$
Tab. 1 shows the results of calculated detection efficiency with different thresholds on SiPM signal and configurations. From the overall perspective of the Tab. 1, it can be observed that as the threshold of the SiPM signal increases, the detection efficiency gradually decreases as expected. From Tab. 1, it can also be observed that $\epsilon_{oneANDother}$ decreases the fastest as the threshold increases. Fig. 11 displayed the energy spectrum under "AND" logic mode. It can be observed that in this mode, there is a significant decrease in the coincidence background.
@@
Tab. 2 shows the rate vs. threshold of the sum of both ends of the PS module, and the coincidence of the two ends when the PS module located at the laboratory on ground without CR monitor. From Tab. 2, it can also be observed that under this "AND" logic mode, the events of backgrounds that coincidentally match sharply decreases. At the same time, it can be found that the rate(63 Hz) corresponding to the threshold of 20 p.e. is close to the muon rate(64 Hz) at sea level.
@@
The CR monitors were also placed at other positions of the PS modules to check the efficiency. When the CR monitors are off to the center of the PS module, the signal strength obtained at the two ends of the PS module will be different. Fig. 12 shows the 2-D plot of the two ends of the PS module when the muon hits around 60 cm away from the center of the PS module. It can be clearly seen that the noise or environmental background can be identified from the signal of the muon. But the signal at one end is significantly larger than the other end, which could generate a difference in the efficiency of the two ends with the same threshold on SiPM signals. The muon signal exhibits elliptical characteristics on the two-dimensional graph, which is due to the asymmetric distance of the CR monitor from the center of PS strip. Combining Fig. 6 and Fig. 6, as well as the two-dimensional graph in Fig. 12, they all indicated that the smaller the MPV of the Landau distribution, the smaller the width of its Landau distribution. When the threshold is too high, even approaching the MPV of one end, the $\epsilon_{oneANDother}$ be determined by the far end (i.e. the end with the smaller signal). Through the measurement results of CR monitor at the outermost edge of PS strip, we found that when the threshold is 15 or 20  p.e., the $\epsilon_{oneANDother}$ is still not affected, almost equivalent to the efficiency at the center. When the threshold is greater than 22  p.e., the $\epsilon_{oneANDother}$ at the edge was approximately 0.3% lower than at the center. By combining with Tab. 2, it is demonstrated that a very good tolerance on the threshold to reach the muon tagging efficiency is possessed by us. Considering the possible efficiency and background under different thresholds, we ultimately decided to use $\epsilon_{module}$ and $\epsilon_{oneANDother}$ as the characterization of overall detection efficiency of PS module.
Fig. 13 shows the relationship between detection efficiency of the sum of the two ends and muon position (CR monitors’ position). The position measurements were made by moving the monitors along the length by a 20 cm step for the 2000mm-PS module (indicated in the figure by dotted lines), or by a 15 cm step for the 1500mm-PS module (indicated in the figure by solid lines). The error bar of the X represents the 6 cm size width of the CR monitor, the Y-coordinate is the detection efficiency of the PS module when the threshold of the sum of the two ends is set to 30 p.e. or 50 p.e., the error bar of the Y-axis is the error in efficiency. As the threshold increases, the detection efficiency corresponding to all positions of the PS module decreases. For a 1500mm-PS module, when the threshold is set to 30 p.e., the detection efficiency of each point is higher than 99.7%. For a 2000mm-PS module, when the threshold is set to 50 p.e., the detection efficiency of each point is higher than 99.6%. Overall, the 1500-mm PS detection efficiency is higher than that of 2000-mm PS. Fig. 13 shows the relationship between detection efficiency of "AND" logic of the two ends and muon position. In this mode, the efficiency of 1500mm-PS is slightly better than that of 2000mm-PS. For all PS module, when the threshold is set to 20 p.e., the detection efficiency of each point is better than 99.3%. The efficiency shows almost no difference along the entire module at 0.1% level.